OpenStudy (anonymous):

How many sides does a polygon have if the measure of each interior angle is twice the measure of each exterior angle?

OpenStudy (anonymous):

interior angle of a polygon= 180(n-2)/n exterior angle =360/n solve

OpenStudy (anonymous):

for a regular polygon!!!!!

OpenStudy (anonymous):

so wait

OpenStudy (anonymous):

each interior angle equals 60 degrees

OpenStudy (anonymous):

and each exterior angle equals 30 degrees

OpenStudy (anonymous):

so there would be 6 sides?

OpenStudy (anonymous):

aye!

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

genius u are

OpenStudy (anonymous):

or was I wrong?

OpenStudy (anonymous):

right